the no of faradays required to deposit 1g  equivalent of Al (at wt=27) from a soln of AlCl3.?

Grade:12

2 Answers

Sunil Kumar FP
9 years ago
1 gm, equivalent of aluminium is equal to 27/3 =9gm
Al+3 +3e- ------------>Al
for 27 gm or 1 mole of aluminium 1F is required
for the decomposition of 9gm faradays of electricty required =F/3
Kushagra Madhukar
3 years ago
Dear student,
Please find the answer to your question.
 
1 gm, equivalent of aluminium is equal to 27/3 = 9gm
Al+3 + 3e- → Al
for 27 gm or 1 mole of aluminium 1F is required
for the decomposition of 9gm faradays of electricty required = F/3
 
Thanks and regards,
Kushagra